OpenStudy (anonymous):

(x+11)/(x+3) • (x^2-9)/(x^2+20x+99) Multiply and simplify

OpenStudy (anonymous):

\[\frac{x + 11}{x + 3} \times \frac{x^{2} - 9}{x^{2} + 20x + 99}\]Can you factor the \(x^{2} - 9\)?

OpenStudy (anonymous):

it would b x-3 right?

OpenStudy (anonymous):

that would be a part of it. What would the other half be?

OpenStudy (anonymous):

x+3 ?

OpenStudy (anonymous):

Yup. \[\frac{x + 11}{x + 3} \times \frac{x^{2} - 9}{x^{2} + 20x + 99} = \frac{x + 11}{x + 3} \times \frac{(x + 3)(x - 3)}{x^{2} + 20x + 99}\]Now can you factor \(x^{2} + 20x + 99\)?

OpenStudy (anonymous):

ummmm no :/

OpenStudy (anonymous):

What are 2 factors of 99 that add up to 20?

OpenStudy (anonymous):

9, 11

OpenStudy (anonymous):

Ok. \[\frac{x + 11}{x + 3} \times \frac{x^{2} - 9}{x^{2} + 20x + 99} = \frac{x + 11}{x + 3} \times \frac{(x + 3)(x - 3)}{(x + 9)(x + 11)} = \]\[\frac{\cancel{x + 11}}{\cancel{x + 3}} \times \frac{\cancel{(x + 3)}(x - 3)}{(x + 9)(\cancel{x + 11})}\]

OpenStudy (anonymous):

what do u do to it now@

OpenStudy (anonymous):

?

OpenStudy (anonymous):

That's literally the answer. Just remove the crossed out parts.

OpenStudy (anonymous):

crossed out parts?

OpenStudy (anonymous):

Yeah. Parts like \[\cancel{x + 3}\]

OpenStudy (anonymous):

ohhhh om thanks

OpenStudy (anonymous):

ok*

OpenStudy (anonymous):

lol. Love how you chose my name ;)

hero (hero):

Yeah, I just typed random letters and it just happened to come out to your name. My bad.